OpenStudy (anonymous):

Write the equation in slope-intercept form.What are the slope and y-intercept? -12x+13y=-10

OpenStudy (the_fizicx99):

Slope intercept form is y =mx + b move the -12x to the right side by adding -12x +13y = -10 +12x +12x divide all the terms by 13 to isolate y 13y= 12x -10 /13 /13 /13 y = 12/13x -0.769 The slope is M in your case its 12/13x The y intercept is the b, in your case thats -0.769

OpenStudy (anonymous):

Thank you so much :)

OpenStudy (the_fizicx99):

b would actually be -0.76923076923076923076923076923077 but I rounded only put the first few digits. You're welcome